OpenStudy (anonymous):

find the sum of three consecutive odd integers if the sum of the first two integers is equal to twenty-four less than four times the third integer.

OpenStudy (anonymous):

\[\text{You're asked to find }n + (n + 2) + (n + 4)\text{ where }n+(n+2)=4(n+4)-24.\\ \ \ \ n+n+2=4n+16-24\\ \ \ \ 2n+2=4n-8\\ \ \ \ -2n=-10\\ \ \ \ n=5\\ \ \ \ n+n+2+n+4=3n+6=15+6=21 \]

OpenStudy (anonymous):

Thank you so much! (: